# Backreaction from quantum fluxes at the Kerr inner horizon

### v1

@@ Abstract
Black holes modeled by the Kerr metric are not semiclassically self-consistent at or below the inner horizon. The renormalized stress-energy tensor (RSET) of a scalar quantum field in the Unruh state has been found to diverge at the Kerr inner horizon, causing the geometry to backreact in a non-trivial way. In an effort to understand this backreaction, here the inner-horizon RSET is computed for the full physically relevant parameter space of black hole spins $a$ and polar angles $\theta$. Then, the backreaction is analyzed using a framework for the dynamical behavior of mass inflation from continued accretion. It is shown that the initial backreaction from the RSET does not evolve the spacetime toward any known regular or extremal configuration, but instead it brings the local interior geometry toward a chaotic, spacelike singularity, classically stable over astrophysical timescales.
@@ I Introduction
In general relativity, a rotating black hole is usually modeled by the Kerr metric, an axisymmetric solution to Einstein’s field equations in an empty spacetime. Within the Kerr black hole, an inner horizon marks the boundary of Cauchy predictability, and beyond it lies an observable singularity, closed timelike curves, and a wormhole to another external universe. But what if the spacetime is not empty? Is it possible for these structures to form in a real gravitational collapse? Put another way, what is the fate of the inner horizon and its traversability if a Kerr black hole is subject to perturbations?
In the case of classical perturbations, the inner horizon is subject to the mass inflation instability [1], an exponential divergence of the locally measured stress-energy due to the presence of both ingoing and outgoing streams of radiation or matter accreted onto the black hole. The backreaction from this classical instability generically results in the geometry’s collapse toward a strong, chaotic, oscillatory, spacelike singularity [2, 3, 4, 5, 6, 7, 8, 9], though if the black hole remains isolated from all external influences aside from its own Price tail, the backreaction could result in a weak, null singularity [10, 11].
In light of the classical divergence of stress-energy near the inner horizon, one may wonder what role quantum effects play in the backreaction as the curvature approaches the Planck scale. If a field in a vacuum state ${|\psi\rangle}$ is canonically quantized over a curved spacetime, its stress-energy contribution to one-loop order can be renormalized and used as a source for semiclassical field equations [12, 13] (throughout, assume geometrized units where $G$=$c$$\hbar$=1):
@@
The computation of this renormalized stress-energy tensor (RSET) is a difficult task for physically realistic vacuum states in even the most symmetric spacetimes. But in recent years, novel computational methods have led to a resurgence of interest, and most relevant to the present analysis, Ref. [14] made use of state subtraction to calculate the mode-summed RSET of a scalar field at the inner horizon of a four-dimensional (4D) Kerr black hole (see [15] for a thorough derivation and Fig. 1 for the parameter space covered by the study). The main conclusion of the study was that the ingoing double-null flux component of the RSET (the $vv$-component, in Eddington-Finkelstein coordinates) is generically nonvanishing, implying that the semiclassical stress-energy will diverge when cast into coordinates that are regular across the inner horizon (as seen, for example, in the local frame of an infalling, outgoing [^1] Infalling (outfalling) here means that the radial component of the observer’s 4-velocity satisfies ${\dot{r}<0}$ (${\dot{r}>0}$), while ingoing (outgoing) refers to a left-moving (right-moving) observer on a traditional Penrose diagram. Ingoing observers are not considered here, since the Unruh state differs substantially from the more realistic Minkowski in-state along that portion of the inner horizon. observer reaching the inner horizon). The implications are much stronger than the classical case—even in the complete absence of external matter or radiation tails, the mere presence of an initially empty quantum field is enough to interrupt the gravitational collapse toward the full Kerr spacetime once the inner horizon is reached.
@@
Figure 1: Parameter space of Kerr spin-to-mass ratios ${a/M}$ and polar angles $\theta$ for which the dominant components of the RSET ${\langle T_{\mu\nu}^{-}\rangle}$ have been calculated for a scalar field at the inner horizon. Cyan plus signs indicate values calculated in Ref. [14] via point splitting, solid blue triangles indicate values calculated in Ref. [14] via state subtraction, and hollow orange circles show the values calculated in the present study via state subtraction.
@@
The divergence of the RSET at the inner horizon implies that the full geometry of the Kerr metric is not semiclassically self-consistent and must be substantially modified by the field’s backreaction. Ref. [14] suggested that the backreaction should depend crucially on the signs of the RSET’s null flux components, based on prior results from the spherically symmetric case [16]: if the RSET’s $vv$-component at the inner horizon is positive (negative), an observer approaching the inner horizon would experience an abrupt contraction (expansion). Since this component was generically found to change signs for different latitudes along the inner horizon and for different black hole spins, the result to first order is a chaotic inner-horizon singularity with local patches of abrupt contraction or expansion.
@@
A more comprehensive analysis of the semiclassical backreaction in rotating black holes was recently carried out in the context of the Kerr-de Sitter spacetime, confirming the postulations of Ref. [14] that the RSET’s $vv$-component dominates the evolution [17]. But if the RSET’s $vv$-component changes sign across different latitudes, this would imply that some portions of the inner horizon would necessarily remain unscathed by any quantum null fluxes. Could a finely-tuned observer evade the semiclassical singularity at the inner horizon? Not quite—Ref. [17] found that even in these regions, a subdominant divergence in the RSET’s $v\varphi$-component causes the local patch of geometry to experience a diverging amount of relative twist.
@@
The goal of the present study is to provide independent validation of the prior studies’ numerical work, extend the parameter space of known RSET values (see Fig. 1), and understand as much as can be possibly gleaned about the Kerr metric’s semiclassical evolution out of equilibrium. There is a hope that the semiclassical gravity framework may be sufficient to describe the (meta)stable endpoint of evolution from a gravitational collapse toward a black hole [18]—though the black hole may evaporate on super-cosmological timescales, the mass-inflation and semiclassical instabilities operate on light-crossing timescales and may push the black hole toward a stable configuration with extremal horizons, a regular core, a wormhole throat, or no horizon at all (such as a gravastar). However, here evidence is given against the formation of such objects from Kerr-like initial conditions. Instead, an analysis of the backreaction using the mass-inflationary framework of Refs. [4, 5, 6] suggests that the local geometry near the would-be inner horizon should evolve in a Belinskii-Khalatnikov-Lifschitz (BKL)-like fashion [19] toward a strong, spacelike singularity.
@@
The derivation of the key details in the calculation of the RSET is presented in Sec. II, the numerical results of the calculations are presented in Sec. III, and the analysis of the backreaction is given in Sec. IV. In this analysis, it should be noted here that one cannot actually conclude definitively that a spacelike singularity is the generic outcome of the semiclassical instability near the inner horizon. The backreaction of the Kerr RSET is only valid while the geometry can still be well-approximated by the vacuum Kerr metric, and as soon as the spacetime begins to evolve away from equilibrium, a new, non-vacuum RSET would need to be found. Most prior studies remain silent on the search for a path forward in our understanding other than an admission to the necessity of a full theory of quantum gravity. But here, another path is suggested: the local inner-horizon geometry under some mild assumptions becomes elegantly simple during the initial stages of mass inflation, and it thus may be possible in the future to compute the RSET in this simplified spacetime to determine the full semiclassical evolution of the near-inner-horizon geometry until the curvature reaches the Planck scale.
@@ II Renormalization of the stress-energy tensor
In order to calculate the Unruh-state RSET at the inner horizon, Ref. [14] constructed a bare mode sum composed of the appropriate differential operator acting on the Hadamard two-point function. Since this sum is formally divergent, the summand was then subtracted from that of another bare mode sum equipped with a vacuum state known to lead to a vanishing RSET at the inner horizon, in order to yield a finite RSET. This latter vacuum state, constructed from a time-reversed, negative-mass Kerr spacetime, may appear unconventional but nonetheless agrees with the Unruh-state RSET obtained from a “comparison” state subtraction [17] as well as from a traditional point splitting approach [14] (at least for two different values of the black hole spin parameter on the polar axis; see Fig. 1).
Computing the two-point function at the inner horizon for a field initialized at the past null boundaries necessitates solving a relativistic scattering problem. Although the Unruh-state modes at the past horizon are not eigenmodes of the radial Teukolsky equation, they can nonetheless be Fourier-decomposed into an orthonormal set of eigenmodes $R^{\text{in}}_{\omega\ell m}$ (originating from past null infinity) and $R^{\text{up}}_{\omega\ell m}$ (originating from the past horizon) so that the two-point function and therefore the RSET at the inner horizon can be expressed in terms of numerically attainable 1D scattering coefficients [15]. These coefficients are calculated using the Mano-Suzuki-Takasugi (MST) method [26, 27], using a version of the Black Hole Perturbation Toolkit (BHPT) [28] adapted by the author; see Appendix A of Ref. [25] for details. In the notation of MST [^2] The translation from the notation of MST, BHPT, and the present work to that of Ref. [14] can be accomplished by setting $\rho^{\text{up}}_{\omega\ell m}={(1-a^{2})^{-2i\omega_{+}(1-a^{2})^{-1/2}}C^{\text{ref}}_{\text{ext}}/C^{\text{inc}}_{\text{ext}}}$ and $B_{\omega\ell m}={(1-a^{2})^{im/a}(r_{-}/r_{+})^{1/2}B^{\text{trans}}_{\text{int}}/B^{\text{trans}}_{\text{ext}}}$, with the remaining coefficients obtained from the Wronskian conditions of Ref. [25]., these scattering coefficients satisfy the asymptotic conditions
@@
where ${\omega_{\pm}\equiv\omega-m\Omega_{r_{\pm}}}$ and the tortoise coordinate $r^{*}$ is chosen to be
@@ III Numerical results
Figure 2: Spectrum of ${\langle T_{vv}^{-}\rangle}$ at the north or south pole (${\theta=0^{\circ}}$) for a Kerr black hole with ${a/M=0.2}$. The sum over the $\ell$- and $m$-modes in the inset panel yields each point in main panel (note that only the ${m=0}$ modes contribute when ${\theta=0^{\circ}}$), and the integral over the $\omega$-modes in the main panel yields the total RSET ${\langle T_{vv}^{-}\rangle}$ (i.e., the point at ${a/M=0.2}$ in Fig. 3). Positive (negative) values on the log plots are given by the red (blue) points. The sampling and integration is performed using the adaptive quadrature algorithm described in the text.
Figure 3: Computed values of the $vv$- and $uu$-components of the RSET at the Kerr inner horizon along the axis of rotation for different values of the black hole spin-to-mass ratio ${a/M}$. Positive (negative) values on the log plot are given by the red (blue) points. Solid curves interpolate between the numerically-computed points, and the dashed curve shows a 1/$a^{2}$ fit at low spins.
@@
Then, the algorithm proceeds by (i) computing four ${\langle T_{\mu\nu}^{-}\rangle_{(\omega)}}$ points subdivided evenly in a closed frequency domain, (ii) interpolating between these points with a cubic B-spline, (iii) computing a fifth sample point in the center of the domain, (iv) calculating the difference between this sample point and the center point of the spline, and (v) if this difference lies above a set error threshold, the domain is divided evenly in half and the algorithm is repeated from step (i) for each new domain. This approach leads to a speedup of up to a factor of 10 in the total computational runtime compared to a fixed linear integration method and additionally allows for more precise error control.
@@
Adaptive quadrature approaches perform most poorly in regimes where the integrand cannot be well-approximated by polynomial functions, which for the integrands of Eqs. (6a)–(6c) occurs in the exponential decay regime at high frequencies (notice in Fig. 2 the slight difference at ${\omega/M\sim 50}$ between the interpolating function and the dashed exponential fit). In these cases, two options were tested, both yielding similar convergent results: either the splines can be computed over log-frequency space and the domain extended until convergence is reached, or enough points can be sampled so that the remaining portion of the integrand can be fitted to an exponentially decaying function and the integral extrapolated to infinity.
@@
Two main sources of error are accounted for in the numerical calculations of the RSET performed here. The first is truncation error, which is minimized in the $\ell$- and $m$-sums by cutting off the sum only when the next term returns zero with the specified numerical precision, and which is controlled in the $\omega$-integrals by the degree of confidence in the exponential decay fit that is integrated to infinity. The second source of error is the global discretization error from the numerical integration scheme, which is controlled by specifying an error bound in the algorithm described above. Accounting for both of these sources of uncertainty, the points in Figs. 24 are computed to a high enough precision that their error bars are smaller than the points themselves in all but a couple edge cases.
@@
First, consider the inner-horizon RSET at the north or south pole, where ${\theta=0^{\circ}}$. There only the modes with ${m=0}$ contribute to the RSET, so that ${\langle T_{v\varphi}^{-}\rangle}$ vanishes and the calculation of the other null components simplify considerably. The two double-null flux components of the RSET are plotted as a function of the black hole’s angular momentum $a$ in Figs. 3 and 4. Fig. 3 shows computed values for spins from ${a/M=0.15}$ to ${a/M\approx 0.997}$, yielding excellent quantitative agreement with Fig. 2 of Ref. [14] and qualitative agreement with the Kerr-de Sitter case in Fig. 4 of Ref. [17] (both of which only reach a minimum of ${a/M=0.55}$). The $vv$-component of the RSET is negative and vanishingly small at high spins, and as the rotation of the black hole slows, this component increases until reaching zero at ${a/M\approx 0.862}$ and continuing to increase exponentially in ${a/M}$. However, at slow enough spins, ${\langle T_{vv}^{-}\rangle}$ once again changes signs.
@@
Both ${\langle T_{vv}^{-}\rangle}$ and ${\langle T_{uu}^{-}\rangle}$ are expected to diverge as ${a\to 0}$, since in that limit the inner horizon coincides with the ${r=0}$ singularity. In particular, the quantity ${\langle T_{uu}^{-}\rangle-\langle T_{vv}^{-}\rangle}$, related to the Hawking outflux per surface area ${4\pi r_{-}^{2}}$ [16], should diverge as $a^{-2}$ (compare the dashed curve in Fig. 3) owing to the factor of $r_{-}$ in the denominator of Eq. (II). An even stronger divergence is expected to be present in ${\langle T_{vv}^{-}\rangle}$ as ${a\to 0}$, since the scattering coefficients ${|B^{\text{ref}}_{\text{int}}|^{2}}$ and ${|B^{\text{trans}}_{\text{int}}|^{2}}$ from Eq. (4a) both diverge as $a^{-2}$. However, the exact nature of the low-spin divergence of ${\langle T_{vv}^{-}\rangle}$ is not easily found, since the spectrum includes non-trivial contributions from successively higher frequencies $\omega$ as the spin $a$ decreases, so that for ${a/M<0.15}$, ${\langle T_{vv}^{-}\rangle}$ could be either positive or negative (or even vanish in a highly fine-tuned case).
@@
While the double-null components of the RSET are expected to diverge at the inner horizon as ${a/M\to 0}$, they will vanish as ${a/M\to 1}$, as shown in Fig. 4 (compare Fig. 3 in the Supplemental Materials of Ref. [14]). In particular, through an analysis of the asymptotic behavior of the near-extremal scattering coefficients, it can be shown that in terms of a near-extremal spin parameter ${\epsilon\equiv\sqrt{1-(a/M)^{2}}}$, the $vv$-component of the RSET will vanish as ${\langle T_{vv}^{-}\rangle\to\epsilon^{4}/(7680\pi^{2})}$, while the $uu$-component will vanish as ${\langle T_{uu}^{-}\rangle\to\epsilon^{5}/(960\pi^{2})}$.
@@
Though not plotted here, the $v\varphi$-component of the RSET has also been computed and is generically non-zero throughout the 2D parameter space, except for at least one 1D zero-valued contour (just as ${\langle T^{-}_{vv}\rangle}$ in Fig. 5 contains two [^3] Or three, if you count the $\theta$-parametrized contour at ${a/M=1}$, though in this work the extremal case is not treated. zero-valued contours given by the white regions). Though the divergence associated with ${\langle T_{v\varphi}^{-}\rangle}$ is subdominant compared to ${\langle T_{vv}^{-}\rangle}$, the presence of a sign flip in ${\langle T_{vv}^{-}\rangle}$ implies that ${\langle T_{v\varphi}^{-}\rangle}$ will dominate the backreaction for at least one latitude, leading to a divergent twisting of the geometry separating regions of local expansion and contraction [17].
In conclusion, the renormalized stress-energy of a scalar quantum field in the Unruh state has been calculated for most of the physically relevant parameter space in the Kerr spacetime at the inner horizon. The results indicate that in a locally inertial frame reaching the inner horizon, the field’s flux generically diverges, and thus that the Kerr metric is not semiclassically self-consistent at or beyond the inner horizon. To understand how the geometry backreacts to this semiclassical instability, one must analyze the field equations, as shown in the next section.

### v2

@@ Abstract
Black holes modeled by the Kerr metric are not semiclassically self-consistent at or below the inner horizon. The renormalized stress-energy tensor (RSET) of a scalar quantum field in the Unruh state has been found to diverge at the Kerr inner horizon [1], causing the geometry to backreact in a nontrivial way. In an effort to understand this backreaction, here the inner-horizon RSET is computed for the full physically relevant parameter space of black hole spins $a$ and polar angles $\theta$. Then, the backreaction is analyzed using a framework for the dynamical behavior of mass inflation from continued accretion. It is shown that the initial backreaction from the RSET does not evolve the spacetime toward any known regular or extremal configuration, but instead it brings the local interior geometry toward a chaotic, spacelike singularity, classically stable over astrophysical timescales.
@@ I Introduction
In general relativity, a rotating black hole is often modeled by the Kerr metric, an axisymmetric solution to Einstein’s field equations in an empty spacetime. Within the Kerr black hole, an inner horizon marks the boundary of Cauchy predictability, and beyond it lies an observable singularity, closed timelike curves, and a wormhole to another external universe. But what if the spacetime is not empty? Is it possible for these structures to form in a real gravitational collapse? Put another way, what is the fate of the inner horizon and its traversability if a Kerr black hole is subject to perturbations?
In the case of classical perturbations, the inner horizon is subject to the mass inflation instability [2], an exponential divergence of the locally measured stress energy due to the presence of both ingoing and outgoing streams of radiation or matter accreted onto the black hole. The backreaction from this classical instability generically results in the geometry’s collapse toward a strong, chaotic, oscillatory, spacelike singularity [3, 4, 5, 6, 7, 8, 9, 10], though if the black hole remains isolated from all external influences aside from its own Price tail, the backreaction could result in a weak, null singularity [11, 12].
In light of the classical divergence of stress energy near the inner horizon, one may wonder what role quantum effects play in the backreaction as the curvature approaches the Planck scale. If a field in a vacuum state ${|\psi\rangle}$ is canonically quantized over a curved spacetime, its stress-energy contribution to one-loop order can be renormalized and used as a source for semiclassical field equations [13, 14] (throughout, assume geometrized units where $G$=$c$$\hbar$=1):
@@
The computation of this renormalized stress-energy tensor (RSET) is a difficult task for physically realistic vacuum states in even the most symmetric spacetimes. But in recent years, novel computational methods have led to a resurgence of interest, and—most relevant to the present analysis—Ref. [1] made use of state subtraction to calculate the mode-summed RSET of a scalar field at the inner horizon of a four-dimensional (4D) Kerr black hole (see [15] for a thorough derivation and Fig. 1 for the parameter space covered by the study). The main conclusion of the study was that the ingoing double-null flux component of the RSET (the $vv$-component, in Eddington-Finkelstein coordinates) is generically nonvanishing, implying that the semiclassical stress energy will diverge when cast into coordinates that are regular across the inner horizon (as seen, for example, in the local frame of an infalling, outgoing [^1] Infalling (outfalling) here means that the radial component of the observer’s 4-velocity satisfies ${\dot{r}<0}$ (${\dot{r}>0}$), while ingoing (outgoing) refers to a left-moving (right-moving) observer on a traditional Penrose diagram. Ingoing observers are not considered here, since the Unruh state differs substantially from the more realistic Minkowski in-state along that portion of the inner horizon. observer reaching the inner horizon). The implications are much stronger than in the classical case—even in the complete absence of external matter or radiation tails, the mere presence of an initially empty quantum field is enough to interrupt the gravitational collapse toward the full Kerr spacetime once the inner horizon is reached.
@@
Figure 1: Parameter space of Kerr spin-to-mass ratios ${a/M}$ and polar angles $\theta$ for which the dominant components of the RSET ${\langle T_{\mu\nu}^{-}\rangle}$ have been calculated for a scalar field at the inner horizon. Cyan plus signs indicate values calculated in Ref. [1] via point-splitting, solid blue triangles indicate values calculated in Ref. [1] via state subtraction, and hollow orange circles show the values calculated in the present study via state subtraction.
@@
The divergence of the RSET at the inner horizon implies that the full geometry of the Kerr metric is not semiclassically self-consistent and must be substantially modified by the field’s backreaction. Reference [1] suggested that the backreaction should depend crucially on the signs of the RSET’s null flux components, based on prior results from the spherically symmetric case [16]: if the RSET’s $vv$-component at the inner horizon is positive (negative), an observer approaching the inner horizon would experience an abrupt contraction (expansion). Since this component was generically found to change signs for different latitudes along the inner horizon and for different black hole spins, the result to first order is a chaotic inner-horizon singularity with local patches of abrupt contraction or expansion.
@@
A more comprehensive analysis of the semiclassical backreaction in rotating black holes was recently carried out in the context of the Kerr–de Sitter spacetime, confirming the postulations of Ref. [1] that the RSET’s $vv$-component dominates the evolution [17]. But if the RSET’s $vv$-component changes sign across different latitudes, this would imply that some portions of the inner horizon would necessarily remain unscathed by any quantum null fluxes. Could a finely tuned observer evade the semiclassical singularity at the inner horizon? Not quite—Ref. [17] found that even in these regions, a subdominant divergence in the RSET’s $v\varphi$-component causes the local patch of geometry to experience a diverging amount of relative twist.
@@
The goal of the present study is to provide independent validation of the prior studies’ numerical work, extend the parameter space of known RSET values (see Fig. 1), and understand as much as can be possibly gleaned about the Kerr metric’s semiclassical evolution out of equilibrium. There is a hope that the semiclassical gravity framework may be sufficient to describe the (meta)stable end point of evolution from a gravitational collapse toward a black hole [18]—though the black hole may evaporate on supercosmological timescales, the mass-inflation and semiclassical instabilities operate on light-crossing timescales and may push the black hole toward a stable configuration with extremal horizons, a regular core, a wormhole throat, or no horizon at all (such as a gravastar). However, here evidence is given against the formation of such objects from Kerr-like initial conditions. Instead, an analysis of the backreaction using the mass-inflationary framework of Refs. [5, 6, 7] suggests that the local geometry near the would-be inner horizon should evolve in a Belinskii-Khalatnikov-Lifschitz (BKL)-like fashion [19] toward a strong, spacelike singularity.
@@
The derivation of the key details in the calculation of the RSET is presented in Sec. II, the numerical results of the calculations are presented in Sec. III, and the analysis of the backreaction is given in Sec. IV. In this analysis, it should be noted here that one cannot actually conclude definitively that a spacelike singularity is the generic outcome of the semiclassical instability near the inner horizon. The backreaction of the Kerr RSET is only valid while the geometry can still be well approximated by the vacuum Kerr metric, and as soon as the spacetime begins to evolve away from equilibrium, a new, nonvacuum RSET would need to be found. Most prior studies remain silent on the search for a path forward in our understanding other than an admission to the necessity of a full theory of quantum gravity. But here, another path is suggested: the local inner-horizon geometry under some mild assumptions becomes elegantly simple during the initial stages of mass inflation, and it thus may be possible in the future to compute the RSET in this simplified spacetime to determine the full semiclassical evolution of the near-inner-horizon geometry until the curvature reaches the Planck scale.
@@ II Renormalization of the stress-energy tensor
In order to calculate the Unruh-state RSET at the inner horizon, Ref. [1] constructed a bare mode sum composed of the appropriate differential operator acting on the Hadamard two-point function. Since this sum is formally divergent, the summand was then subtracted from that of another bare mode sum equipped with a vacuum state known to lead to a vanishing RSET at the inner horizon, in order to yield a finite RSET. This latter vacuum state, constructed from a time-reversed, negative-mass Kerr spacetime, may appear unconventional but nonetheless agrees with the Unruh-state RSET obtained from a “comparison” state subtraction [17] as well as from a traditional point-splitting approach [1] (at least for two different values of the black hole spin parameter on the polar axis; see Fig. 1).
Computing the two-point function at the inner horizon for a field initialized at the past null boundaries necessitates solving a relativistic scattering problem. Although the Unruh-state modes at the past horizon are not eigenmodes of the radial Teukolsky equation, they can nonetheless be Fourier-decomposed into an orthonormal set of eigenmodes $R^{\text{in}}_{\omega\ell m}$ (originating from past null infinity) and $R^{\text{up}}_{\omega\ell m}$ (originating from the past horizon), so that the two-point function and therefore the RSET at the inner horizon can be expressed in terms of numerically attainable 1D scattering coefficients [15]. These coefficients are calculated using the Mano-Suzuki-Takasugi (MST) method [26, 27], using a version of the Black Hole Perturbation Toolkit (BHPT) [28] adapted by the author; see Appendix A of Ref. [25] for details. In the notation [^2] The translation from the notation of MST, BHPT, and the present work to that of Ref. [1] can be accomplished by setting $\rho^{\text{up}}_{\omega\ell m}={(1-a^{2})^{-2i\omega_{+}(1-a^{2})^{-1/2}}C^{\text{ref}}_{\text{ext}}/C^{\text{inc}}_{\text{ext}}}$ and $B_{\omega\ell m}={(1-a^{2})^{im/a}(r_{-}/r_{+})^{1/2}B^{\text{trans}}_{\text{int}}/B^{\text{trans}}_{\text{ext}}}$, with the remaining coefficients obtained from the Wronskian conditions of Ref. [25]. of MST, these scattering coefficients satisfy the asymptotic conditions
@@
where ${\omega_{\pm}\equiv\omega-m\Omega_{r}(r_{\pm})}$, and the tortoise coordinate $r^{*}$ is chosen to be
@@ III Numerical results
Figure 2: Spectrum of ${\langle T_{vv}^{-}\rangle}$ at the north or south pole (${\theta=0^{\circ}}$) for a Kerr black hole with ${a/M=0.2}$. The sum over the $\ell$- and $m$-modes in the inset panel yields each point in the main panel (note that only the ${m=0}$ modes contribute when ${\theta=0^{\circ}}$), and the integral over the $\omega$-modes in the main panel yields the total RSET ${\langle T_{vv}^{-}\rangle}$ (i.e., the point at ${a/M=0.2}$ in Fig. 3). Positive (negative) values on the log plots are given by the red (blue) points. The sampling and integration is performed using the adaptive quadrature algorithm described in the text.
@@
Then, the algorithm proceeds by (i) computing four ${\langle T_{\mu\nu}^{-}\rangle_{(\omega)}}$ points subdivided evenly in a closed frequency domain, (ii) interpolating between these points with a cubic B-spline, (iii) computing a fifth sample point in the center of the domain, (iv) calculating the difference between this sample point and the center point of the spline, and (v) if this difference lies above a set error threshold, dividing the domain evenly in half and repeating the algorithm from step (i) for each new domain. This approach leads to a speedup of up to a factor of 10 in the total computational runtime compared to a fixed linear integration method and additionally allows for more precise error control.
@@
Adaptive quadrature approaches perform most poorly in regimes where the integrand cannot be well approximated by polynomial functions, which for the integrands of Eqs. (6a)–(6c) occurs in the exponential decay regime at high frequencies (notice in Fig. 2 the slight difference at ${\omega/M\sim 50}$ between the interpolating function and the dashed exponential fit). In these cases, two options were tested, both yielding similar convergent results: either the splines can be computed over log-frequency space and the domain extended until convergence is reached, or enough points can be sampled so that the remaining portion of the integrand can be fitted to an exponentially decaying function and the integral extrapolated to infinity.
@@
Two main sources of error are accounted for in the numerical calculations of the RSET performed here. The first is truncation error, which is minimized in the $\ell$- and $m$-sums by cutting off the sum only when the next term returns zero with the specified numerical precision, and which is controlled in the $\omega$-integrals by the degree of confidence in the exponential decay fit that is integrated to infinity. The second source of error is the global discretization error from the numerical integration scheme, which is controlled by specifying an error bound in the algorithm described above. Accounting for both of these sources of uncertainty, the points in Figs. 24 are computed to a high enough precision that their error bars are smaller than the points themselves in all but a couple of edge cases.
@@
First, consider the inner-horizon RSET at the north or south pole, where ${\theta=0^{\circ}}$. There, only the modes with ${m=0}$ contribute to the RSET, so that ${\langle T_{v\varphi}^{-}\rangle}$ vanishes and the calculation of the other null components simplifies considerably. The two double-null flux components of the RSET are plotted as a function of the black hole’s angular momentum $a$ in Figs. 3 and 4. Figure 3 shows computed values for spins from ${a/M=0.15}$ to ${a/M\approx 0.997}$, yielding excellent quantitative agreement with Fig. 2 of Ref. [1] and qualitative agreement with the Kerr–de Sitter case in Fig. 4 of Ref. [17] (both of which only reach a minimum of ${a/M=0.55}$). The $vv$-component of the RSET is negative and vanishingly small at high spins, and as the rotation of the black hole slows, this component increases until reaching zero at ${a/M\approx 0.862}$ and continuing to increase exponentially in ${a/M}$. However, at slow enough spins, ${\langle T_{vv}^{-}\rangle}$ once again changes signs.
@@
Both ${\langle T_{vv}^{-}\rangle}$ and ${\langle T_{uu}^{-}\rangle}$ are expected to diverge as ${a\to 0}$, since in that limit the inner horizon coincides with the ${r=0}$ singularity. In particular, the quantity ${\langle T_{uu}^{-}\rangle-\langle T_{vv}^{-}\rangle}$, related to the Hawking outflux per surface area ${4\pi r_{-}^{2}}$ [16], should diverge as $a^{-2}$ (compare the dashed curve in Fig. 3) owing to the factor of $r_{-}$ in the denominator of Eq. (II). An even stronger divergence is expected to be present in ${\langle T_{vv}^{-}\rangle}$ as ${a\to 0}$, since the scattering coefficients ${|B^{\text{ref}}_{\text{int}}|^{2}}$ and ${|B^{\text{trans}}_{\text{int}}|^{2}}$ from Eq. (4a) both diverge as $a^{-2}$. However, the exact nature of the low-spin divergence of ${\langle T_{vv}^{-}\rangle}$ is not easily found, since the spectrum includes nontrivial contributions from successively higher frequencies $\omega$ as the spin $a$ decreases, so that for ${a/M<0.15}$, ${\langle T_{vv}^{-}\rangle}$ could be either positive or negative (or even vanish in a highly fine-tuned case).
Figure 3: Computed values of the $vv$- and $uu$-components of the RSET at the Kerr inner horizon along the axis of rotation for different values of the black hole spin-to-mass ratio ${a/M}$. Positive (negative) values on the log plot are given by the red (blue) points. Solid curves interpolate between the numerically computed points, and the dashed curve shows a 1/$a^{2}$ fit at low spins.
@@
While the double-null components of the RSET are expected to diverge at the inner horizon as ${a/M\to 0}$, they will vanish as ${a/M\to 1}$, as shown in Fig. 4 (compare Fig. 3 in the Supplemental Material of Ref. [1]). In particular, through an analysis of the asymptotic behavior of the near-extremal scattering coefficients, it can be shown that in terms of a near-extremal spin parameter ${\epsilon\equiv\sqrt{1-(a/M)^{2}}}$, the $vv$-component of the RSET will vanish as ${\langle T_{vv}^{-}\rangle\to\epsilon^{4}/(7680\pi^{2})}$, while the $uu$-component will vanish as ${\langle T_{uu}^{-}\rangle\to\epsilon^{5}/(960\pi^{2})}$.
@@
Though not plotted here, the $v\varphi$-component of the RSET has also been computed and is generically nonzero throughout the 2D parameter space, except for at least one 1D zero-valued contour (just as ${\langle T^{-}_{vv}\rangle}$ in Fig. 5 contains two [^3] Or three, if you count the $\theta$-parametrized contour at ${a/M=1}$, though in this work the extremal case is not treated. zero-valued contours given by the white regions). Though the divergence associated with ${\langle T_{v\varphi}^{-}\rangle}$ is subdominant compared to ${\langle T_{vv}^{-}\rangle}$, the presence of a sign flip in ${\langle T_{vv}^{-}\rangle}$ implies that ${\langle T_{v\varphi}^{-}\rangle}$ will dominate the backreaction for at least one latitude, leading to a divergent twisting of the geometry separating regions of local expansion and contraction [17].
In conclusion, the renormalized stress energy of a scalar quantum field in the Unruh state has been calculated for most of the physically relevant parameter space in the Kerr spacetime at the inner horizon. The results indicate that in a locally inertial frame reaching the inner horizon, the field’s flux generically diverges, and thus that the Kerr metric is not semiclassically self-consistent at or beyond the inner horizon. To understand how the geometry backreacts to this semiclassical instability, one must analyze the field equations, as shown in the next section.